# Wasserstein Flow Matching: Generative modeling over families of distributions

### v1

@@ E.1 Spatial Transcriptomics
In our manuscript, we applied WFM and BW-FM on several spatial transcriptomics datasets, encompassing a variety of technologies and tissue contexts. From a 254-gene MERFISH atlas of the motor cortex (Zhang et al.,, 2021), we focus on niches of microglia cells. We compress gene-expression profiles down to their $16$ principal components (PC) and aggregate all the cells around each microglia within an $80$ micron radius, yielding on average $26.6$ cells per niche. We then calculated the gene-expression PC mean and covariance within each environment to produce Gaussians for BW-FM. Generated Gaussians align with real data and span for the full cortical depth of the microglia niches (Figure S2). In Figure 3, we predict the environment composition by cell type for generated Gaussians via nearest-neighbour regression in BW space using real data as supervision, demonstrating congruence between the two across cortical depth.
@@
In a complementary approach, WFM is applied directly on gene-expression based point-clouds of niches, and does not require the Gaussian representation. Uniquely suited for high-dimensional data, we apply WFM on seqFISH assay of embryogenesis (Lohoff et al.,, 2022) and a XENIUM experiment of melanoma metastasis to the brain (Haviv et al., 2024b, ). In both dataset, we select the $k=8$ physical nearest neighbours of every cell, and aggregate their first $16$ PCs to produce environment point-clouds.
From the seqFISH dataset, we concentrated on the gut-tube region, which is divided into spatially segregated, gastrulating organs. Applied unconditionally, WFM generated niches match the distribution of the real data based on EMD and CD 1-nearest-neighbour accuracy (Table 2). We then assessed WFM’s capability to comprehend the relationship between cell state and environment and tasked it with conditional generation based on organ label. Based on OT distances estimated via Wormhole embeddings (Haviv et al., 2024a, ), point-clouds from WFM recapitulated true organ environment.The label accuracy for WFM-generated data was 78.86%, which was nearly identical to the test-set real data accuracy of 79.59%.
Appendix F 2D & 3D point-clouds
Figure S4: Cities Dataset. Gaussians representing the $100$ most populous cities in the continental US. The data was obtained from Bennett, (2010) via OSMnx (Boeing,, 2017). The mean parameter is the longitude and latitude coordinate of each city and the covariance is the 2nd moment approximation of their metro area.
The ShapeNet dataset consists of 3D point-clouds of $55$ different classes, each one comprised of $15,000$ points. Emulating the benchmarking effort in (Wu et al.,, 2023), we apply WFM to generate $n=1000$ sized examples from the plane, car and chair classes. At each gradient descent step, we sample $64$ point-clouds from the training set for each class, and randomly select $n=1000$ points from each. To evaluate generation quality, we synthesize point-clouds to much the size of the test set, and calculate the real or generated $1-NN$ accuracy based on EMD and CD metrics.
ModelNet has $40$ classes of point-clouds, with $2048$ points in each. Conditioned on class label, WFM is trained to generate $n=1000$ sized point-clouds here too. In this setting, the noise measure is the standard normal and we did not use multi-sample matching. According to nearest-neighbour classification from OT preserving Wormhole embeddings, generated samples match their class with an accuracy of $77.66\%$, approaching the $79.98\%$ purity of test set samples from real-data.
The MNIST dataset is a widely used collection of handwritten digits, consisting of 28x28 pixel grayscale images of the numbers 0 through 9. EMNIST (Extended MNIST) is an expansion of MNIST that includes handwritten letters as well as digits. To convert samples from these datasets into point-clouds, we threshold each image and extract the coordinates of the above-threshold pixels. This produces a cohort of point-clouds of variables sizes, as each image contains a different number of relevant pixel. We apply the entropic OT map (see Appendix A) based WFM to synthesize point-clouds of the digit $4$ and letter $a$. Despite the data heterogeneity, WFM produces realistic examples (Figure 5), while capturing the data distribution (Table 2). We again stress that this is a unique feature of WFM, lacking from any previous point-cloud generation algorithm

### v2

@@ F.1 Spatial Transcriptomics
In our manuscript, we applied WFM and BW-FM on several spatial transcriptomics datasets, encompassing a variety of technologies and tissue contexts. From the 351-gene seqFISH embryogensis dataset, (Lohoff et al., 2022), we focus on niches of gut-tube cells. We compress gene-expression profiles down to their $16$ principal components (PC) and aggregate all the cells around each gut cell within an $80$ micron radius, yielding on average $29$ cells per niche. We then calculated the gene-expression PC mean and covariance within each environment to produce Gaussians for BW-FM. Generated Gaussians align with real data ad delineate budding organs into their appropriate regions (see Figure S2).
Figure S3: Conditional BW-FM applied to single-cell RNA sequencing data of immune response to COVID-19. Large scale single-cell atlases are commonly grouped into highly dedicated clusters called MetaCells (Persad et al., 2023). In this application, BW-FM is conditioned on cell state and trained to generate means and covariances of gene expression, focusing on the top 32 principal components, derived from aggregated cells. The model achieves high-quality sample generation, as evidenced by a label accuracy of 93.13%.
In a complementary approach, WFM is applied directly on gene-expression based point-clouds of niches, representing a realization of the continuous distribution of their environments. Uniquely suited for high-dimensional data, we apply WFM on a MERFISH atlas of the motor cortex XENIUM assay of melanoma metastasis to the brain (Haviv et al., 2024b). In both dataset, we select the $k=16$ physical nearest neighbours of every cell, and aggregate expression profiles based on their first $16$ PCs.
We concentrated on Sst interneurons in the motor cortex datas, which are divided into spatially segregated, phenotypically distinct subtypes. Applied unconditionally, WFM generated niches match the distribution of the real data based on EMD and CD 1-nearest-neighbour accuracy (see Table 4). We then assessed WFM’s capability to comprehend the relationship between cell state and environment and tasked it with conditional generation based on subtype label. Based on OT distances estimated via Wormhole embeddings (Haviv et al., 2024a), distributions generated by WFM recapitulated true organ environment. The label accuracy for WFM-generated data was 63.86%, which was nearly identical to the test-set real data accuracy of 62.19%.
Figure S4: Cities Dataset. Gaussians representing the $100$ most populous cities in the continental US. The data was obtained from (Bennett, 2010) via OSMnx (Boeing, 2017). The mean parameter is the longitude and latitude coordinate of each city and the covariance is the 2nd moment approximation of their metro area.
Appendix G 2D & 3D shapes
The ShapeNet dataset consists of 3D shapes of $55$ different classes, Emulating the benchmarking effort in (Wu et al., 2023a), we apply WFM to generate $n=2048$ sized examples from the plane, car and chair classes. At each gradient descent step, we sample $64$ shapes from the training set for each class, and randomly select $n=2048$ points from each. To evaluate generation quality, we synthesize point-clouds to much the size of the test set, and calculate the real or generated $1-NN$ accuracy based on EMD and CD metrics.
@@
ModelNet has $40$ classes of shapes, each consisting of $n=2048$ particles. Conditioned on class label, WFM is trained to generate $n=1000$ sized point-clouds here too. In this setting, the noise measure is the standard normal and we did not use multi-sample matching. According to nearest-neighbour classification from OT preserving Wormhole embeddings, generated samples match their class with an accuracy of $77.66\%$, approaching the $79.98\%$ purity of test set samples from real-data.
The MNIST dataset is a widely used collection of handwritten digits, consisting of 28x28 pixel grayscale images of the numbers 0 through 9. EMNIST (Extended MNIST) is an expansion of MNIST that includes handwritten letters as well as digits. To convert samples from these datasets into distributions, we threshold each image and extract the coordinates of the above-threshold pixels. This produces a cohort of point-clouds of variables sizes, as each image contains a different number of relevant pixels. We apply the entropic OT map (see Appendix A) based WFM to synthesize distributions of the digit $4$ and letter $a$. Despite the data heterogeneity, WFM produces realistic examples (see Figure 4), while capturing the data distribution (see Table 4).